# Approaching Rate-Distortion Limits in Neural Compression with Lattice Transform Coding

### v1

@@ 4.2 Entropy Models
Specifically, the analysis transform $g_{a}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{y}}$ maps each vector of the block to its latent variable, $\bm{y}_{i}=g_{a}(\bm{x}_{i})$, which has $d_{y}$ dimensions. This produces a latent i.i.d. vector sequence $\bm{y}_{1},\dots,\bm{y}_{n}$. Then, $d_{y}$ separate LTC models $\{c_{a}^{(i)},c_{s}^{(i)},p_{\hat{\bm{y}}}^{(i)}\}_{i=1}^{d_{y}}$ are applied to across the block, where $c_{a}^{(i)},c_{s}^{(i)}$ are the analysis/synthesis transforms of LTC. The $i$-th LTC model is applied to the latent slice $[\bm{y}_{1,i},\dots,\bm{y}_{n,i}]^{\top}$, which has i.i.d. scalar entries. The resulting reconstructed latents $\hat{\bm{y}}_{i}$’s are passed to the synthesis transform $g_{s}:\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d}$, such that $\hat{\bm{x}}_{i}=g_{s}(\hat{\bm{y}}_{i})$.
@@ 5 Results
We analyze the three settings of interest: (i) i.i.d. scalar sequences, (ii) general vector sources, and (iii) i.i.d. vector sequences. In addition, we provide an ablation study demonstrating the effects of each component in LTC (lattice choice, quantization training scheme, density model, transform choice).
5.1 Experimental Setup
The objective function used is the rate-distortion lagrangian (1), with STE used for the distortion term. For the rate term, we found varying results for dithered quantization (13) and STE (11) depending on the source; these are specified for each experiment individually. Additional implementation details can be found in Appendix A.
Lattices.
We choose several commonly used lattices that are known to be the best in their dimension (Agrell and Allen, 2023): hexagonal, $D^{*}_{n}$, $E_{8}$ and $\Lambda_{24}$ (Leech) lattices. These lattices all have efficient algorithms to solve the closest vector problem (CVP), which computes $Q_{\Lambda}(\bm{y})$. The algorithms used are referenced in Sec. 2.
Transforms.
For the scalar i.i.d. sequences and general vector sources, we choose the transforms $g_{a}$ and $g_{s}$ to be MLPs with two hidden layers, a hidden dimension of 100, and softplus nonlinearities after each layer (except the last). For the vector i.i.d. sources, where the lattice transform coding is applied in the latent space, we apply additional companding MLPs of the same architecture for the $c_{a}^{(i)}$ and $c_{s}^{(i)}$ transforms.
Entropy Models.
We use the RealNVP (Dinh et al., 2017) normalizing flow for the density model $p_{\bm{y}}$, with 5 coupling layers and isotropic Gaussian prior. The rates plotted (unless mentioned otherwise) are given by the cross-entropy upper bound (14) of the density model. We use 4096 samples for the Monte-Carlo estimate of the PMF in (11) for all lattices. This was sufficient for most sources and lattices; see Appendix. B for a discussion on this topic.
5.2 Compression of Scalar i.i.d. Sequences
@@ 5.1 Experimental Setup
We consider i.i.d. sequences of several sources, including Gaussian with $L_{2}$ distortion, Laplacian with $L_{1}$ distortion, and a 1-dimensional marginal of the Banana source (Ballé et al., 2021) with $L_{2}$ distortion. The 1-d Banana marginal is used to test a skewed distribution, since both Gaussian and Laplacian sources are symmetric. In this setting, $n$ is the dimension of both $\bm{x}$ and the latent space $\bm{y}$, and also represents the block-length with regards to the rate-distortion function. Here, one-shot coding of $\bm{x}$ is equivalent to block coding of the scalar marginal source. We first discuss the 2-dimensional Gaussian case, which provides visualization of what LTC achieves before expanding to more general dimensions. We found that using STE for the rate term performed better in this setting (see Sec. 5.3).
2-dimensions.
(a) Original space.
(b) Latent space.
@@
Figure 6: LTC quantization regions for $n=2$, using the hexagonal lattice $A_{2}$, on i.i.d. Gaussians.
@@
In 2-dimensions, we choose $Q_{\Lambda}$ to be the hexagonal lattice, which is known to be optimal (Conway and Sloane, 1999), and use the NF density model. As shown in Figs. 5(a), LTC produces quantization regions in the original source space that closely match that of optimal ECVQ in Fig. 3(b). The latent space (Fig. 5(b)) effectively uses the hexagonal lattice, compared to NTC’s use of the integer lattice (Fig. 13). The analysis transform $g_{a}$ determines which source realizations $\bm{x}$ get quantized to a specific lattice point in the latent space (colored regions); the synthesis transform $g_{s}$ maps the lattice points to their corresponding reconstructions in the original space (black dots). Regarding rate-distortion performance, the rate-redundancy matches that of ECVQ in 2-dimensions.
$n$-dimensions.
Here, we choose $Q_{\Lambda}$ to be the best-known lattices in $n$-dimensions, including $A_{2}$ (hexagonal), $D_{4}^{*}$, $E_{8}$, and $\Lambda_{24}$. As $n$ grows, ECVQ performance with approach the rate-distortion function of the source, which for the Gaussian (squared error distortion) is given by $R(D)=\frac{1}{2}\log\left(\frac{\sigma^{2}}{D}\right),0\leq D\leq\sigma^{2}$ and the Laplacian (absolute error distortion) is given by $R(D)=-\log(\lambda D),0\leq D\leq 1/\lambda$. as provided by Cover and Thomas (2006); Equitz and Cover (1991). For the Banana 1-d marginal, we use Blahut-Arimoto (Blahut, 1972; Arimoto, 1972) to estimate $R(D)$ numerically since it is unknown in closed-form.
Figure 7: Rate-redundancy of Gaussian i.i.d. sequences.
(a) Laplace ($L_{1}$ distortion).
(b) Banana-1D ($L_{2}$ distortion).
Figure 8: Rate-distortion performance of scalar i.i.d. sequences.
We plot $\mathcal{E}_{\mathsf{c}}(D)$ against $D$ and $R$. On the Gaussian source, we see that in Fig. 7, LTC performs very close to optimal ECVQ in dimensions $n=2,4,8$ using the hexagonal, $D^{*}_{n}$, and $E_{8}$ lattice respectively. For larger $n$, ECVQ is too costly to compute, but we see that LTC with the Leech lattice for $n=24$ continues the trend, further reducing the rate-redundancy. We observe similar performance gains for the Laplacian in Fig. 14; we do not show the rate-redundancy for Banana-1d since its rate-distortion function does not have a closed-form.
Since each of the lattice quantizers used have efficient CVP algorithms, LTC succeeds in larger dimensions and rates, whereas ECVQ could not even be run due to its complexity that is exponential in dimension and rate. Note that ECVQ performance provides a lower bound on $\mathcal{E}_{\mathsf{c}}(D)$ for any fixed $n$, which indicates that LTC is successfully leveraging space-filling efficiency in higher dimensions.
A standard rate-distortion figure is shown in Figs. 1,8, where LTC outperforms NTC by a constant rate reduction for a fixed SNR, in all three sources. This provides numerical evidence that a companding performance gap corresponding to the lattice’s packing efficiency such as (5) also holds for non-asymptotic rates.
@@ 5.3 Ablation Study
(a) LTC with $A_{n}$ and $D^{*}_{n}$ lattices.
(b) Dither vs. STE for rate term.
Figure 9: Ablation study, i.i.d. Gaussian sequences. Solid lines are ECVQ, following legend in Fig. 7.
We analyze (i) the lattice choice, (ii) the training objective used, and (iii) the effect of density models on the entropy estimate of $\hat{\bm{y}}$. For the lattice choice, only using the best-known lattices in each dimension (in terms of NSM) yielded near-ECVQ performance. If sub-optimal lattices are used, this was not the case. Fig. 8(a) shows the $A_{n}$ and $D^{*}_{n}$ lattices used for $n=8$; they perform about the same as ECVQ for $n=4$, but still yield coding gain over NTC.
For the training objective used, we compare using dithered quantization with STE for the rate term (i.e., using likelihoods from (13) and (11) respectively). Shown in Fig. 8(b), for low rates, the reported cross-entropies from dithering are much larger than that of STE, but converge to the STE performance as rate increases. One possible reason is that the convolved density in (13) over-smooths the true density of $\bm{y}$ in the low-rate regime, leading to sub-optimality.
5.4 One-Shot Compression of Vector Sources
In general, $\bm{x}$ is a vector source with non-i.i.d. entries. NTC and LTC function as one-shot codes for $\bm{x}$, and we use $d$ to refer to the dimensionality of $\bm{x}$, and $d_{y}$ the latent dimensionality. This setting encompasses one-shot compression of the Sawbridge (Wagner and Ballé, 2021) as well as the full 2-d Banana source (Ballé et al., 2021). These works demonstrate that NTC already recovers optimal one-shot coding on these sources, so there is no gain possible over NTC for one-shot codes. Instead, we consider real-world sources such as the Physics and Speech datasets (Yang and Mandt, 2022), where $d=16$ and $33$ respectively. It is unknown if NTC is one-shot optimal on these sources.
Figure 10: One-shot compression of the Speech source.
For NTC, $\bm{x}$ is mapped to a $d$-dimensional latent space. For LTC, we map $\bm{x}$ to a latent space with dimension of the lattice used. For these sources, we found training instabilities with STE, so dithering is used instead. At higher rates, we also observed stabler training with the factorized density model compared to the normalizing flow; more details in appendix. For Speech, we observe coding gain for rates above $\approx 7$ bits with the $D_{4}^{*}$ and $E_{8}$ lattices, shown in Fig. 10. Shown in Fig. 17, coding gain is observed on Physics for LTC in different rate regimes. LTC with the hexagonal lattice performs best at lower rates, but the performance drops off as rate increases. At higher rates, LTC with 4 latent dimensions performs the best with $D_{4}^{*}$ lattice.
5.5 Block Compression of i.i.d. Vector Sources
For a general vector source $\bm{x}$, gains over the optimal one-shot coding performance can be achieved by coding i.i.d. blocks of $\bm{x}$. We apply BLTC to the Sawbridge and Banana sources, for which NTC was shown to be one-shot optimal. For these i.i.d. vector sources, we again use STE for the BLTC models.
(a) Sawbridge.
(b) Banana.
Figure 11: Performance of BLTC on i.i.d. sequences of vector sources, where $n$ is the block-length.
For the Sawbridge, the optimal one-shot performance is given by $H(D)$ (Wagner and Ballé, 2021, Thm. 2), and $R(D)$ is estimated using NERD (Lei et al., 2022). Shown in Figs. 10(a), 15, we verify that BLTC bridges the gap between $H(D)$ and $R(D)$. In this case, BLTC uses a latent dimension $d_{y}=1$, and the nominal dimension is $d=1024$. For the Banana source, we estimate $R(D)$ using Blahut-Arimoto, and again, LTC bridges the gap between NTC performance and $R(D)$. Here, BLTC uses a latent dimension of $d_{y}=2$ and $d=2$. Thus, we verify on two sources that BLTC can outperform optimal one-shot codes due to the use of block coding. Finally, for real-world data, on the Physics source we additionally see coding gains with BLTC, in Fig. 18. For Physics, optimal one-shot coding is unknown (improved one-shot coding was achieved via LTC in the previous section); however, block compression is able to outperform NTC.
6 Conclusion
We show that NTC is sub-optimal on i.i.d. sequences due to its use of the integer lattice. By leveraging better lattices in higher dimensions, LTC compresses i.i.d. sequences nearly optimally, improves one-shot coding performance, and enables the design of i.i.d. vector block compressors.
7 Acknowledgements
We sincerely thank Lucas Theis for discussions which helped develop many ideas in this work.

### v2

@@ 4.3 From One-Shot to Block Coding
LTC can be directly applied to any vector source $\bm{x}\in\mathbb{R}^{d}$. In this case, LTC operates as a one-shot code for compressing $\bm{x}$. In Sec 5, we discuss the improvements that LTC provides compared to NTC in the one-shot sense. Rate-distortion theory informs us that to go beyond one-shot performance and approach asymptotic optimality (given by $R_{\bm{x}}(D)$), we need to encode i.i.d. copies $\bm{x}_{1},\dots,\bm{x}_{n}\sim p_{\bm{x}}$ simultaneously. This is highly relevant in settings such as dataset compression, when we wish not just to compress a single sample, but an entire dataset where samples can be assumed i.i.d. Note that $n$ remains the block-length, but $d$ is now the dimension of $\bm{x}$.
To realize such a compressor, we propose Block-LTC (BLTC), which applies LTC across the i.i.d. latent variables of $\bm{x}$, shown in Fig. 4. Specifically, the analysis transform $g_{a}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{y}}$ maps each vector of the block to its latent variable, $\bm{y}_{i}=g_{a}(\bm{x}_{i})$, which has $d_{y}$ dimensions. This produces a latent i.i.d. vector sequence $\bm{y}_{1},\dots,\bm{y}_{n}$. Then, $d_{y}$ separate LTC models $\{c_{a}^{(i)},c_{s}^{(i)},p_{\hat{\bm{y}}}^{(i)}\}_{i=1}^{d_{y}}$ are applied to across the block, where $c_{a}^{(i)},c_{s}^{(i)}$ are the analysis/synthesis transforms of LTC. The $i$-th LTC model is applied to the latent slice $[\bm{y}_{1,i},\dots,\bm{y}_{n,i}]^{\top}$, which has i.i.d. scalar entries. The resulting reconstructed latents $\hat{\bm{y}}_{i}$’s are passed to the synthesis transform $g_{s}:\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d}$, such that $\hat{\bm{x}}_{i}=g_{s}(\hat{\bm{y}}_{i})$. We use STE to train the BLTC models. The BLTC architecture first transforms the i.i.d. vectors into i.i.d. latent vectors $\bm{y}_{i}$; the space-packing ability of LTC compresses the i.i.d. latent vectors.
@@ 5 Results
We first discuss LTC on synthetic sources, before moving to real-world sources. Then, we demonstrate how LTC performs with nested lattices, and finally present results on block coding using BLTC. In addition, we provide an ablation study demonstrating the effects of various components in LTC.
Experimental setup. For LTC’s rate term, we found varying results for dithering (6) and STE (5) depending on the source; these are specified for each experiment individually. Additional implementation details can be found in Appendix A. We choose lattices best in their dimension (Agrell & Allen, 2023): $A_{2}$ (hexagonal), $D^{*}_{4}$, $E_{8}$ and $\Lambda_{24}$ (Leech) lattices. These lattices all have efficient algorithms to solve the closest vector problem (CVP), which computes $Q_{\Lambda}(\bm{y})$. For the Euclidean vector sources, we choose the transforms $g_{a}$ and $g_{s}$ to be MLPs with two hidden layers, a hidden dimension of 100, and softplus nonlinearities after each layer (except the last). For tensor-valued sources, we use standard source-specific transforms (e.g., CNNs for images). Unless mentioned otherwise, we use the RealNVP (Dinh et al., 2017) normalizing flow (NF) for the density model $p_{\bm{y}}$ (used for variable-rate and fixed-rate LTC), with 5 coupling layers and isotropic Gaussian prior. The rates plotted (unless mentioned otherwise) are given by the cross-entropy of the density model. We use 4096 samples for the Monte-Carlo estimate of the PMF in (5) and (8) for all lattices. This was sufficient for most sources and lattices; see Appendix. A.2. In most experiments, we compare with ECVQ as well as the rate-distortion function $R(D)$ of the source. When not known in closed-form, $R(D)$ is estimated using Blahut-Arimoto (BA) (Blahut, 1972; Arimoto, 1972) for 1-D and 2-D sources, and NERD (Lei et al., 2022) for higher-dimensional sources.
5.1 LTC as a One-Shot Code
@@
We first consider 2-D isotropic Gaussians and the 2-D Banana source, where we can easily visualize the quantization regions. Then, we consider $n$-length i.i.d. sequences of several sources, including Gaussian with $L_{2}$ distortion, Laplacian with $L_{1}$ distortion, and a 1-dimensional marginal of the Banana source (Ballé et al., 2021) with $L_{2}$ distortion. The 1-d Banana marginal is used to test a skewed distribution, since both Gaussian and Laplacian sources are symmetric. In this setting, $n$ is the dimension of both $\bm{x}$ and the latent space $\bm{y}$. We found that using STE for the rate term worked better for the i.i.d. sources, whereas dithering was more stable for the Banana source.
@@
Figure 5: LTC ($A_{2}$ lattice) quantization regions. Left: 2-D Gaussians. Right: Banana.
Figure 6: LTC on synthetic sources. Left: Gaussians. Right: Banana.
@@
2-D Gaussian and Banana. In 2-dimensions, we choose $Q_{\Lambda}$ to be the hexagonal lattice, which is known to be optimal (Conway & Sloane, 1999), and use the NF density model. As shown in Figs. 6, LTC produces quantization regions in the original source space that closely match that of optimal ECVQ in Fig. 3. The latent space (Fig. 17(b)) effectively uses the hexagonal lattice, compared to NTC’s use of the integer lattice (Fig. 17(a)). The analysis transform $g_{a}$ determines which source realizations $\bm{x}$ get quantized to a specific lattice point in the latent space (colored regions); the synthesis transform $g_{s}$ maps the lattice points to their corresponding reconstructions in the original space (black dots). Shown in Fig. 6, the LTC rate-distortion performance matches that of ECVQ for the Gaussian. On the Banana, NTC and LTC achieve ECVQ at low rates; at larger rates, LTC improves on NTC but does not fully match ECVQ; see Fig. 6. Inspecting the quantization regions (Fig. 21, 21, 21) reveals why. At low rates, all methods quantize along a 1-D nonlinear component of the source, resulting in the same quantization regions no matter the tessellating structure. At higher rates, 2 dimensions are used for quantization, where the integer lattice of NTC becomes sub-optimal compared to the hexagonal regions of ECVQ/LTC. We see that the quantizer used solely determines LTC’s quantization regions; whereas the transforms’ purpose is to warp the latent quantizer onto the source statistics.
$n$-length i.i.d. Sequences. Here, we choose $Q_{\Lambda}$ to be the best-known lattices in $n$-dimensions, including $A_{2}$ (hexagonal), $D_{4}^{*}$, $E_{8}$, and $\Lambda_{24}$. As $n$ grows, ECVQ performance will approach the rate-distortion function of the source, which for the Gaussian (squared error distortion) is given by $R(D)=\frac{1}{2}\log\left(\frac{\sigma^{2}}{D}\right),0\leq D\leq\sigma^{2}$ and the Laplacian (absolute error distortion) is given by $R(D)=-\log(\lambda D),0\leq D\leq 1/\lambda$ (Cover & Thomas, 2006; Equitz & Cover, 1991). On Gaussians, we see that in Fig. 6, LTC performs very close to optimal ECVQ in dimensions $n=2,4,8$ using the hexagonal, $D^{*}_{n}$, and $E_{8}$ lattice respectively. For larger $n$, ECVQ is too costly to compute, but we see that LTC with the Leech lattice for $n=24$ continues the trend, further improving the rate-distortion performance and approaching $R(D)$. We observe similar performance gains for the Laplacian and Banana-1D in Fig. 16. Since each of the lattice quantizers used have efficient CVP algorithms, LTC succeeds in larger dimensions and rates, whereas ECVQ could not even be run due to its exponential complexity. This result provides evidence that companding (in the form of LTC) is near-optimal in terms of VQ performance, even if it may not exactly implement VQ (Gersho, 1979).
Real-World Sources. We now consider sources $\bm{x}$ drawn from real-world data. We consider the Physics and Speech datasets (Yang & Mandt, 2022), of dimension $d=16$ and $33$ respectively, and the Kodak image dataset (Franzen, ), which has variable dimension depending on image size.
Figure 7: LTC on real-world sources. Left: Speech. Middle two: Physics. Right: Kodak.
For Physics and Speech, we map $\bm{x}$ to a latent space with dimension $d_{y}$ using MLP-based transforms. For these sources, we found training instabilities with STE, so dithering is used instead. At higher rates, we also observed stabler training with the factorized density model compared to the normalizing flow. For Speech, we observe coding gain for rates above $\approx 7$ bits with the $D_{4}^{*}$ and $E_{8}$ lattices, shown in Fig. 7. Coding gain is observed on Physics for LTC in different rate regimes. LTC with the hexagonal lattice performs best at lower rates, but the performance drops off as rate increases. At higher rates, LTC with 4 latent dimensions performs the best with $D_{4}^{*}$ lattice. These rates are measured in bits-per-sample, and not normalized by dimension. Generally speaking, different rate regimes may require a different number of latent dimensions to be used for NTC and LTC. If the number of latent dimensions used by the transforms does not match the lattice quantizer dimension, this can result in suboptimal performance as the tessellation in source space may not be optimal. This is likely why LTC performance “peaks” for different latent dimensions.
For large-scale images, we focus on the NTC architecture in Cheng et al. (2020), and use product $E_{8}$ and Leech $\Lambda_{24}$ lattices along the channel dimension of the latent. We use dither during training and STE at test time, and train on the Vimeo-90k (Xue et al., 2019) dataset. We use the identical hyperparameter settings for transforms and entropy models, corresponding to the first 4 quality levels for the “cheng2020-attn” model in CompressAI (Bégaint et al., 2020). Shown in Fig. 7, LTC with product $E_{8}$ and product Leech achieves a -5.274% and -16.708% BD-rate gain respectively over NTC when evaluated on Kodak. In Sec. B, we compare all methods with Cheng2020-NTC and VTM as anchor in the BD-rate table (Tab. 1). While NVTC achieves a -2.919 BD-rate gain over VTM, Cheng2020 with Leech is significantly better, with a -11.895% BD-rate gain. Note that NVTC and Cheng2020 models differ significantly in terms of the transforms and entropy models, so multiple factors (e.g., transforms and quantization) may contribute to the achieved performance. Additionally, NVTC applies vector quantization to vectors of size 16, meaning NVTC is able to leverage the packing efficiency of 16-dimensional space compared to the $E_{8}$ and Leech with 8 and 24 respectively. To compare with LVQAC, we additionally plot the performance using the $D_{n}^{*}$ lattice, which achieves a -2.337% BD-rate gain over Cheng2020-NTC. This compares solely the effect of quantization scheme, and not other components of LVQAC (Zhang & Wu, 2023). This result, along with how the $E_{8}$ and $\Lambda_{24}$ perform relative to each other, further supports evidence that the rate-distortion performance is significantly determined by the quantization choice in the latent space, and that the NTC model is able to leverage the enhanced packing efficiency of higher-dimensional lattices. We see that simply adding lattice quantization to the latent space of Cheng2020 has brought its performance to approach the ResNet-VAE rate-distortion upper bound $R_{U}(D)$ from Yang & Mandt (2022) on Kodak. To compare just the speed of SQ vs LQ, we run a forward pass through the transforms, quantization, and entropy model. When run on single-core CPU, Cheng2020-NTC takes 3.241 seconds per image averaged over Kodak, whereas Cheng2020-LTC-E8 takes 4.002 seconds per image. Additional comments on complexity can be found in Sec. A, B, with full runtime comparison in Tab 2.
As NTC and LTC operate here as one-shot compressors of $\bm{x}$, these results demonstrate that the gap to $R(D)$ is partially due to NTC being sub-optimal as a one-shot code. Fully approaching $R(D)$, however, requires encoding across multiple source realizations, which we discuss in Sec. 5.2.
Nested Lattice LTC. Here, we demonstrate the ability of fixed-rate LTC with nested lattices to provide coding gain with lattice dimension, with a much lower complexity when coding to bits. On i.i.d. Gaussians (Fig. 9), we demonstrate LTC with NLQ in the latent space. We use self-similar nested lattices for the integer, $E_{8}$ and $\Lambda_{24}$ lattices, and set the nesting ratio $\Gamma=5,7,9$ to sweep the rate. The performance improves as the lattice dimension grows, and approaches $R(D)$ as before, but does not perform as well as variable-rate LTC for a particular dimension, as expected. We note that LTC with NLQ has encoding complexity linear in $n$ and is able to outperform NTC.
5.2 Block Coding
We evaluate BLTC in the block coding setup (i.e., compressing multiple samples from dataset simultaneously) on the Sawbridge, Banana, and Physics sources. For the Sawbridge, the optimal VQ (i.e., one-shot coding) performance is known in closed form as $H(D)$ (Wagner & Ballé, 2021, Thm. 2). In this case, BLTC uses a latent dimension $d_{y}=1$, and the nominal dimension is $d=1024$. In Fig. 9, we verify that BLTC bridges the gap between $H(D)$ and $R(D)$. Note that going beyond $H(D)$ is impossible without compressing multiple realizations at time; this verifies the ability of BLTC to leverage space-packing of the Sawbridge’s latent source. For the Banana and Physics sources, the optimal one-shot performance is unknown, but we see that BLTC improves upon NTC and LTC, which are both one-shot codes. Thus, we show BLTC’s ability to approach $R(D)$.
@@ 5.3 Ablation Study
We first show how it is necessary for a good lattice to be used to achieve optimal performance. Only using the best-known lattices in each dimension yielded near-ECVQ performance. If sub-optimal lattices are used, this was not the case. Fig. 15(c) shows the $A_{n}$ and $D^{*}_{n}$ lattices used for $n=8$; they perform about the same as ECVQ for $n=4$, but still yield coding gain over NTC. This implies that LVQAC, which uses $D_{n}^{*}$ is sub-optimal compared to other lattices. In the appendix (Sec. B.1), we provide several additional ablations further revealing the interplay of LTC components.
6 Discussion and Limitations
Figure 8: Block coding. Left: Sawbridge. Center: Banana. Right: Physics.
Figure 9: LTC-NLQ.
In this work, we use the best known lattices up to dimension 24. Going to even larger dimensions will improve performance further; designing efficient high dimensional lattices with low complexity CVP solvers is an active area of research building on advances in modern coding theory (e.g. low density parity codes and polar codes).
Acknowledgements
This work was supported by The Institute for Learning-enabled Optimization at Scale (TILOS), under award number NSF-CCF-2112665. The work of Hamed Hassani was further supported by NSF CAREER award CIF-1943064, and the work of Eric Lei was further supported by a NSF Graduate Research Fellowship.